HETAV PATEL

Grade 11,

P,Q,R,S are four vector of equal magnitude. If P + Q - R= 0 angle between P and Q is theta 1. If P + Q - S = 0 angle between P and S is theta2. the ratio of theta 1 and theta 2 is

P,Q,R,S are four vector of equal magnitude. If P + Q - R= 0 angle between P and Q is theta 1. If P + Q - S = 0 angle between P and S is theta2. the ratio of theta 1 and theta 2 is

Grade:12

1 Answers

Vikas TU
7 years ago
Write P +Q – R as
P + Q = R
take modulus both sides and square it.
|P|^2 + |Q|^2 + 2P.Q = |R|^2
Now all have same magnitude,
P.Q = -|R|^2/2 ….......................(1)
cos(thetha1)  = P.Q/|R|^2 from eqn. (1)
put  it in costhetha1
costhetha1 = -1/2
thetha1 = 120 degree.
Similarly U ccan find now thetha2.